Let E and F be two events of an experiment with sample space S. Suppose P(e)=.6 P(F)=.4 and P( E & F)=.2 compute:
a) P(E U F)
b) P(E')
c) P(F')
d) P (E' & F)

Answer by stanbon(75887)

Let E and F be two events of an experiment with sample space S. Suppose
P(E)=.6 P(F)=.4 and P( E & F)=.2 compute:
a) P(E U F) = P(E)+P(F)-P(E & F) = 0.6+0.4-0.2 = 0.8
-------------------------
b) P(E') = 1-P(E) = 0.4
----------------------------
c) P(F') = 1-P(F) = 0.6
---------------------
d) P (E' & F) = P(E')+P(F)-P(E'U F)
-------------
Cheers,
Stan H.
-----------